OpenStudy (anonymous):

What is the midpoint of (–2, 0) and (4, –2)?

OpenStudy (mathlegend):

\[\frac{ x1 + x2 }{ 2 } \frac{ y1 + y2 }{ 2}\]

OpenStudy (anonymous):

(-2 + 4)/2 = 1 (0 + -2)/2 = -2 So (1,-2) right?

hartnn (hartnn):

(0-2)/2 = ??

OpenStudy (anonymous):

Oh ok so (1, -1) right?

hartnn (hartnn):

yup :)

OpenStudy (mathlegend):

Great job @jazy

OpenStudy (anonymous):

Thanks @MathLegend & @hartnn (:

OpenStudy (anonymous):

How do I find the length of one side of triangle(on a coordinate grid)?

OpenStudy (mathlegend):

I think you can use the distance formula

hartnn (hartnn):

u have both endpoints , right ?

OpenStudy (anonymous):

They don't give any measures, this is the picture and they ask for the length of line BC

OpenStudy (mathlegend):

or maybe you can find the slope of each side?

OpenStudy (mathlegend):

but I'm thinking distance formula

hartnn (hartnn):

see carefully, u have both end-points !

hartnn (hartnn):

and yes, distance formula it is.

OpenStudy (anonymous):

yeah i just found the distance formula in my notes sorry!!!\[d=\sqrt{(x2 - x1) + (y2 - y1)}\]

OpenStudy (anonymous):

And thanks again! (:

hartnn (hartnn):

no problem,welcome :)